OpenStudy (anonymous):

solve: sin^2theta-cos^2theta= 1+costheta

OpenStudy (anonymous):

Hint: Use the relationship sin^2theta + cos^2theta = 1 to replace the sin^2theta term in your equation. Then you will have an equation with only cos's.(no sin's)

OpenStudy (anonymous):

cos^2theta - sin^2theta= cos(2theta i thought?

OpenStudy (anonymous):

No, don't use that one. Use \[\sin ^{2}\theta + \cos ^{2}\theta = 1\]

OpenStudy (anonymous):

solve that for sin^2(theta) and you get sin^2(theta) = 1 - cos^2(theta)

OpenStudy (anonymous):

Plug that into your original equation for sin^2(theta). Next step : get 0 on 1 side and figure out which values of theta make the other side = 0

OpenStudy (anonymous):

that makes so much more sense!!! thanks

OpenStudy (anonymous):

no problem